A firm has short run TC:


TC = 100 + 2 Q + Q2 with MC = 2 + 2 Q


1) Find ATC and AVC functions


2) If P = 25, how much will the firm produce in the SR?


3) If P = 20, how much will the firm produce in the SR?


4) Assuming the firm has the same cost curves in the LR, how much will it produce in the LR?


1
Expert's answer
2021-11-29T15:12:18-0500

TC = 100 + 2Q + Q^2, MC = 2 + 2Q.

1) "ATC = TC\/Q = 100\/Q + 2 + Q."

AVC = VC/Q = 2 + Q.

2) If P = 25, then the firm will produce the next Q in the SR:

P = MC,

2 + 2Q = 25,

Q = 11.5 units.

3) If P = 20, then the firm will produce the next Q in the SR:

P = MC,

2 + 2Q = 20,

Q = 9 units.

4) If the firm has the same cost curves in the LR, then it will produce the next Q in the LR:

MC = ATC,

100/Q + 2 + Q = 2 + 2Q,

Q - 100/Q = 0,

"Q^2 = 100,"

Q = 10 units.